# Spin Relaxation Mechanisms and Nuclear Spin Entanglement of the V Center in hBN

### v1

@@ Abstract
The negatively charged boron vacancy (V) defect in hexagonal boron nitride (hBN) has recently emerged as a promising spin qubit for sensing due to its high-temperature spin control and versatile integration into van der Waals structures. While extensive experiments have explored their coherence properties, much less is known about the spin relaxation time () and its control parameter dependence. In this work, we develop a parameter-free spin dynamics model based on the cluster expansion technique to investigate relaxation mechanisms at low temperature. Our results reveal that the V center constitutes a strongly coupled electron spin-nuclear spin core, which necessitates the inclusion of the coherent dynamics and derived memory effects of the three nearest-neighbor nitrogen nuclear spins. Using this framework, this work closely reproduces the experimentally observed time at G and further predicts the dependence on external magnetic field in the G interval, when the spin relaxation is predominantly driven by electron–nuclear and nuclear–nuclear flip–flop processes mediated by hyperfine and dipolar interactions. This study establishes a reliable and scalable approach for describing relaxation in V centers and offers microscopic insights to support future developments in nuclear-spin-based quantum technologies.
@@ 1 Introduction
In solid-state quantum systems, spin qubits have emerged as a crucial resource for advancing quantum technologies, including quantum sensing[1, 2], quantum memory[3, 4, 5, 6, 7], quantum computing[8], and quantum repeaters[9]. Among various defect-based platforms, the nitrogen-vacancy (NV) center in diamond has demonstrated remarkable potential, owing to its long spin coherence times and the ability to initialize, coherently manipulate, and optically read out individual electron spins at room temperature[10, 11]. Motivated by these achievements, considerable effort has recently focused on identifying and characterizing other optically addressable spin defects in wide-bandgap materials[1, 12, 13, 14]. In particular, point defects in hexagonal boron nitride have attracted growing interest due to their intrinsically high photon out-coupling efficiency[15], excellent compatibility with photonic components[16, 17], long-term photostability[18, 19], bright and spectrally pure single-photon emission[12], as well as various defect choices[20, 21].
@@
Among these, the negatively charged boron vacancy defect has emerged as a candidate for enabling coherent control of an electron spin and nearby nuclear spins in hBN. The V center consists of a missing boron atom coordinated by three nitrogen atoms in a D3h symmetry, forming an electronic spin system with a zero-field splitting (ZFS) of approximately 3.5 GHz [22, 23, 24]. Recent experimental studies have characterized its spin and optical properties [22, 23, 24, 25, 26, 27]. At room temperature, longitudinal relaxation times have been reported in the range of 15–18 s, largely independent of applied magnetic field and host isotope composition [22, 25, 26]. At cryogenic temperatures, extends to 1 ms up to 12.5 ms, depending on the isotope concentration and the magnetic field strength[22, 25]. Recent experiments have demonstrated that the V center remains stable in ultrathin, few-layer hBN flakes, exhibiting robust optically detected magnetic resonance signals and spin polarization, while spin-lattice relaxation appears to depend on the dimensionality of the sample.[28] However, the comprehensive analysis about dependence of on the external magnetic field remains largely unexplored.
@@
In terms of spin coherence, spin echo times () are strongly influenced by the nuclear spin environment, yielding values of 82.1 ns for natural isotopic abundance hBN[29], 46 ns for h11BN, and 62 ns for h10BN[30]. While most early investigations focused on 14N-containing systems, recent work has identified hB15N as a particularly favorable isotopic configuration, offering improved and times due to reduced hyperfine interactions[22, 31, 32].
@@
Despite significant experimental and theoretical progress in understanding decoherence in V center[22, 30, 33, 34, 31, 32], comprehensive theoretical modeling of their spin relaxation dynamics remains limited. In particular, while several computational studies have focused on investigating time in these systems[31, 32, 35], systematic modeling of the time has been comparatively scarce. A thorough understanding of the mechanisms governing is essential, as it ultimately sets an upper bound for the achievable spin coherence time through the fundamental relation [36]. Typically, the relaxation time comprises multiple contributions arising from distinct microscopic processes, including spin–spin and spin–lattice interactions[25, 37, 38]. At room temperature, spin relaxation is predominantly governed by spin-lattice interactions[22, 39]. This can be theoretically captured by ab initio spin dynamics simulations[40, 41]. However, at cryogenic temperatures, where phonon populations are suppressed, spin relaxation is largely dictated by spin flip-flop processes through hyperfine and dipolar interactions with the surrounding nuclear spin bath. Capturing dipolar spin relaxation is particularly challenging, as the approximations used for computing decoherence in cluster correlation expansion methods are not applicable, and many-spin entanglements may need to be taken into account, especially in the dense nuclear spin environment in hBN[42]. To date, no comprehensive model has been constructed to fully capture these low-temperature spin relaxation mechanisms in V centers.
@@
In this work, we therefore develop a spin dynamics model to investigate the relaxation mechanism of V centers in hBN at low temperature and moderate magnetic fields. To accurately account for the interplay of hyperfine interactions, dipolar couplings, and electron-nuclear entanglement effects, we employ a cluster expansion technique combined with an extended Lindbladian formalism[42, 43]. This approach enables us to approximate the time evolution of the diagonal elements of the reduced density matrix of the central spin system without simplifying the underlying spin interactions, while remaining computationally tractable. Using this framework, we first carry out various models and then identify the most effective model capable of accurately capturing spin decoherence. We show that the electron spin-three first-neighbor 15N nuclear spin coupled system must be treated as a single coherent unit, an extended core, in order to reproduce the experimentally observed decay characteristics. We then apply this optimized model to elucidate the relaxation mechanisms responsible for the experimentally observed at G[22]. Finally, we extend our analysis to predict the dependence on the external magnetic field. These results offer valuable microscopic insight into the spin relaxation dynamics of V centers and establish a theoretical foundation for future developments in coherence control, defect engineering, and device optimization for quantum information applications.
@@ 2 Results
where the first term on the right-hand side of the equation describes the electron Zeeman interaction, while the second and third terms describe the zero-field splitting interaction. Here, is Bohr magneton, is an external magnetic field in the direction perpendicular to the hBN, is a factor of an electron, for an electronic spin-1, is the zero-field splitting equal to 3.479 GHz[24], and is the transverse component of the ZFS, which is set to zero in this work.
@@
where is the hyperfine-interaction tensor, calculated based on the method in Ref. [44] to tackle the finite-size effect. As shown in Eq. (4), the , , , and hyperfine components govern the electron–nuclear spin flip-flop processes. To illustrate the strength of these interactions, their values are plotted in Fig. 1(b), highlighting the strongly coupled nature of the V defect system. Fig. 1(b) unravels that rotational invariant perpendicular hyperfine components of the first three nearest nitrogen nuclei are strong due to the Fermi contact contribution. This suggests that relaxation via the electron–nuclear spin flip-flop process in this channel is indispensable and plays a central role in the overall spin dynamics.
To simulate spin relaxation process, our models take into account all major parts of the spin Hamiltonian in Eq. (1) and truncate only the nuclear spin-nuclear spin dipolar couplings to a certain degree depending on the clustering used. The dimension of the given total Hamiltonian depends on the dimension of the Hilbert space of electron and nuclear spins, and the computational demand increases exponentially with the number of nuclear spins, limiting the capability of the exact solution. To resolve this issue, the methodology developed in Ref. [42] is applied by dividing the entire system into smaller clusters of spins utilizing the central spin approximation and cluster expansion to reduce the Hilbert space considered while still maintaining all the aforementioned interactions.
@@
We simulate the dynamics of the system through the density matrix () following the below Master equation with an extended Lindbladian developed by Ref. [42]
We note that represents an extended Lindbladian term introduced into the Master equation to account for spin flip-flop processes induced by environmental spins not included in cluster (diagonal elements in the density matrix), see Ref. [42] for more details. Additionally, it is used to ensure the synchronization of key properties of different instances of the central spin across different clusters, such as expectation values. This correction is essential because the unsynchronization of a central spin in all clusters can lead to an artificial thermal state due to the accumulation of its magnetization. The extended term can be formulated as follows
Here, determines spin flip-flop rates of the central spin induced by environmental spins not incorporated in cluster and is the jump operator describing the spin flip and flop transitions of the central spin between and states.
@@
2.1 Theoretical Prediction.
As detailed in the Methods section, the time is determined from the population of the electron spin until it relaxes to thermal equilibrium. Fig. S1 in Supplementary S1 illustrates the energy-level splitting among the three relevant states and suggests that the fields in the vicinity of the ground-state level anticrossing (GSLAC) yield particularly narrow splitting, as the GSLAC is the point at which two ground-state sublevels become nearly degenerate. This implies that spin relaxation near GSLAC is faster, which can minimize computational expense. We therefore adopt a fixed field at 800 G to evaluate the performance of various computational models for predicting .
Model 1: Two-Spin Clusters
@@
Fig. 3(a) depicts the evolution of the V electron spin initialized in the ms = 0 state. The result suggests that immediately after initialization, the electron spin polarization is rapidly transferred to the 15N nuclear spin in this model. Subsequently, while the ms = 0 population relaxes toward the thermal equilibrium value of 1/3, as expected for the equal occupancy of the three spin sublevels (m), the relaxation pattern deviates from the exponential curve[45, 46] and exhibits a stretched exponential decay. This is confirmed by the exponent value of 0.37. Stretched exponents often arise when the relaxation is governed by multiple, independent decay mechanisms with different rates. While the distribution of the perpendicular hyperfine components, comprising strongly, moderately, and weakly coupled spins, see Fig. 1(b), could suggest such a multi-channel relaxation mechanism, experimental results do not support this model. As a consequence, we conclude that Model 1 fails to capture relevant correlation effects of the bath that may lead to different decay mechanisms.
Model 2 - Three/Four-Spin Clusters
@@
Unlike Model 1, this model incorporates dipolar interactions within each cluster, which may additionally contribute to spin dynamics behavior. In addition, Model 2 accounts for up to 4 spin correlation effects, thus, it significantly improves on Model 1 and can describe more complex dynamics. Here, 47 nitrogen clusters and 48 boron clusters have been used, describing a 238-spin model in total. Note that the number of clusters has been tested as shown in Supplementary S3.
@@
Model 3 - Three/Four-Spin Clusters with Additional Dephasing
Before moving toward another clusterization, we investigate the effect of docoherence of the electron spin on time. The cluster Hamiltonian in this model remains the same as in Model 2, and the number of clusters is also identical. However, in this model, we incorporate the effects of spin dephasing by introducing a dephasing time () of 200 ns [22] into the off-diagonal terms of the Lindbladian in Eq. (7). Note that the variation of values has been tested in Supplementary S4.
@@
Model 4 - Six-Spin Clusters
@@
To manage computational complexity, we apply the same reduced density matrix approach as in the previous two models. Furthermore, a total of 28 clusters are considered in this model, accounting for a total of 60 spins. It is important to note that in the model, we focus on nitrogen nuclear spins, and boron nuclei are neglected at first, as they exhibit a weak hyperfine interaction and make a combined boron-nitrogen model computationally unfeasible. Therefore, the computed time obtained from Model 4 should be regarded as an upper bound. Note that the timestep convergence test and the effect of nuclear-spin synchronization have been validated in Supplementary S5.
@@
Model 4 enlarges the central‐spin subsystem to include the V electron spin and its three nearest‐neighbor 15N nuclei, while the bath is augmented by two additional distant 15N nuclei. The initial electron–spin polarization in the ms = 0 sublevel is set to 70%, in agreement with experimental measurements[47, 48]. Under these conditions, the expected exponential relaxation of electron spin polarization appears, see Fig. 3(d). The decay behavior is well described by an exponential function (decay exponent equal to 1). This demonstrates that the previously unaccounted correlation effects arise from insufficient entanglement and dipolar interactions within the extended central spin and also with the bath spins. That is, the spin dynamics are mediated by the spin flip-flop processes. Specifically, referring to Eqs. (3) and (4), nuclear–nuclear flip–flop transitions are driven by the dipolar interaction term whereas electron–nuclear flip–flop transitions originate from the non‐secular hyperfine term . As illustrated in Fig. 1(b), the V electron spin exhibits strong hyperfine coupling to its three nearest nitrogen nuclei, therefore, including these four spins in the central‐spin description is essential to accurately capture the unseparable dynamics of the electron spin and the first neighbor nucelar spins. Simultaneously, the addition of two distant nuclear spins enables an accurate description of nuclear–nuclear flip‐flops via the dipolar interaction and additional correlation effects. Model 4 can be considered as defining the boundary between the central spin and the spin bath at the first-neighbor shell of the V center. Consequently, the reduced bath includes only the moderately and weakly coupled nuclear spins. In this extended central spin–moderately coupled spin bath model, the spin bath can be considered “memoryless”, or at least its correlations have no significant influence on the dynamics of the extended central spin system, which, in turn, gives rise to an exponential decay characteristic of Markovian processes. As a result, this model encompasses all significant relaxation pathways accurately and will be employed for all remaining calculations. A more detailed analysis of the contribution of distant nuclear spins is presented in the following section.

### v2

# Nuclear Spin-Mediated Relaxation Mechanisms of the V Center in hBN

@@ Abstract
The negatively charged boron vacancy (V) defect in hexagonal boron nitride has recently emerged as a promising spin qubit for sensing due to its high-temperature spin control and versatile integration into van der Waals structures. While extensive experiments have explored their coherence properties, much less is known about the spin relaxation time () and its control parameter dependence. In this work, we develop a parameter-free spin dynamics model based on the cluster expansion technique to investigate relaxation mechanisms at low temperature. Our results reveal that the V center constitutes a strongly coupled electron spin-nuclear spin core, which necessitates the inclusion of the coherent dynamics and derived memory effects of the three nearest-neighbor nitrogen nuclear spins. Using this framework, this work closely reproduces the experimentally observed time at G and further predicts the dependence on external magnetic field in the G interval, when the spin relaxation is predominantly driven by electron–nuclear and nuclear–nuclear flip–flop processes mediated by hyperfine and dipolar interactions. This study establishes a reliable and scalable approach for describing relaxation in V centers and offers microscopic insights to support future developments in nuclear-spin-based quantum technologies.
@@ 1 Introduction
In solid-state quantum systems, spin qubits have emerged as a crucial resource for advancing quantum technologies, including quantum sensing[10.1038/s41563-024-01887-z, 10.1038/s41467-021-24725-1], quantum memory[10.1103/PhysRevX.6.021040, 10.1038/nphys2026, 10.1126/science.1220513, 10.1063/5.0188597, 10.1002/adom.202302760], quantum computing[10.1063/5.0007444], and quantum repeaters[10.1038/s41534-022-00637-w]. Among various defect-based platforms, the nitrogen-vacancy (NV) center in diamond has demonstrated remarkable potential, owing to its long spin coherence times and the ability to initialize, coherently manipulate, and optically read out individual electron spins at room temperature[10.1016/j.physrep.2013.02.001, 10.1126/science.1139831]. Motivated by these achievements, considerable effort has recently focused on identifying and characterizing other optically addressable spin defects in wide-bandgap materials[10.1038/s41563-024-01887-z, 10.1021/acsnano.3c08940, 10.1038/s41467-024-51129-8, 10.1021/acs.jpclett.3c01475]. In particular, point defects in hexagonal boron nitride have attracted growing interest due to their intrinsically high photon out-coupling efficiency[10.1038/nnano.2015.242], excellent compatibility with photonic components[10.1088/1361-6463/aa7839, 10.1002/adom.202002218], long-term photostability[10.1021/acsphotonics.8b00127, 10.1016/j.physe.2020.114251], bright and spectrally pure single-photon emission[10.1021/acsnano.3c08940], as well as various defect choices[10.3390/nano12142427, 10.1021/acs.jpcc.4c03404].
@@
Among these, the negatively charged boron vacancy defect has emerged as a candidate for enabling coherent control of an electron spin and nearby nuclear spins in hBN. The V center consists of a missing boron atom coordinated by three nitrogen atoms in a D3h symmetry, forming an electronic spin system with a zero-field splitting (ZFS) of approximately 3.5 GHz [10.1038/s41467-023-44494-3, 10.1038/s41563-020-0619-6, 10.1038/s41524-020-0305-x]. Recent experimental studies have characterized its spin and optical properties [10.1038/s41467-023-44494-3, 10.1038/s41563-020-0619-6, 10.1038/s41524-020-0305-x, 10.1126/sciadv.abf3630, 10.1021/acs.nanolett.1c02495, 10.1038/s41563-022-01329-8]. At room temperature, longitudinal relaxation times have been reported in the range of 15–18 s, largely independent of applied magnetic field and host isotope composition [10.1038/s41467-023-44494-3, 10.1126/sciadv.abf3630, 10.1021/acs.nanolett.1c02495]. At cryogenic temperatures, extends to 1 ms up to 12.5 ms, depending on the isotope concentration and the magnetic field strength[10.1038/s41467-023-44494-3, 10.1126/sciadv.abf3630]. Recent experiments have demonstrated that the V center remains stable in ultrathin, few-layer hBN flakes, exhibiting robust optically detected magnetic resonance signals and spin polarization, while spin-lattice relaxation appears to depend on the dimensionality of the sample.[durand_optically_2023] However, the comprehensive analysis about dependence of on the external magnetic field remains largely unexplored.
@@
In terms of spin coherence, spin echo times () are strongly influenced by the nuclear spin environment, yielding values of 82.1 ns for natural isotopic abundance hBN[10.1038/s41467-022-33399-2], 46 ns for h11BN, and 62 ns for h10BN[10.1038/s41467-022-31743-0]. While most early investigations focused on 14N-containing systems, recent work has identified hB15N as a particularly favorable isotopic configuration, offering improved and times due to reduced hyperfine interactions[10.1038/s41467-023-44494-3, 10.1002/adfm.202511300, 10.1038/s41699-022-00336-2].
@@
Despite significant experimental and theoretical progress in understanding decoherence in V center[10.1038/s41467-023-44494-3, 10.1038/s41467-022-31743-0, rizzato_extending_2023, ramsay_coherence_2023, 10.1002/adfm.202511300, 10.1038/s41699-022-00336-2], comprehensive theoretical modeling of their spin relaxation dynamics remains limited. In particular, while several computational studies have focused on investigating time in these systems[10.1002/adfm.202511300, 10.1038/s41699-022-00336-2, 10.1103/PhysRevB.106.104108], systematic modeling of the time has been comparatively scarce. A thorough understanding of the mechanisms governing is essential, as it ultimately sets an upper bound for the achievable spin coherence time through the fundamental relation [10.1088/0034-4885/80/1/016001]. Typically, the relaxation time comprises multiple contributions arising from distinct microscopic processes, including spin–spin and spin–lattice interactions[10.1126/sciadv.abf3630, 10.1103/PhysRevB.97.094304, 10.1038/s41524-021-00673-8]. At room temperature, spin relaxation is predominantly governed by spin-lattice interactions[10.1038/s41467-023-44494-3, 10.1103/PhysRevLett.108.197601]. This can be theoretically captured by ab initio spin dynamics simulations[10.1038/s41524-023-01082-9, 10.1103/PhysRevB.98.214442]. However, at cryogenic temperatures, where phonon populations are suppressed, spin relaxation is largely dictated by spin flip-flop processes through hyperfine and dipolar interactions with the surrounding nuclear spin bath. Capturing dipolar spin relaxation is particularly challenging, as the approximations used for computing decoherence in cluster correlation expansion methods are not applicable, and many-spin entanglements may need to be taken into account, especially in the dense nuclear spin environment in hBN[10.1103/PhysRevB.101.155203]. To date, no comprehensive model has been constructed to fully capture these low-temperature spin relaxation mechanisms in V centers.
@@
In this work, we therefore develop a spin dynamics model to investigate the relaxation mechanism of V centers in bulk hBN at low temperature and moderate magnetic fields. To accurately account for the interplay of hyperfine interactions, dipolar couplings, and electron-nuclear entanglement effects, we employ a cluster expansion technique combined with an extended Lindbladian formalism[10.1103/PhysRevB.101.155203, xu_ab_2024]. This approach enables us to approximate the time evolution of the diagonal elements of the reduced density matrix of the central spin system without simplifying the underlying spin interactions, while remaining computationally tractable. Using this framework, we first carry out various models and then identify the most effective model capable of accurately capturing spin decoherence. We show that the electron spin-three first-neighbor 15N nuclear spin coupled system must be treated as a single coherent unit, an extended core, in order to reproduce the experimentally observed decay characteristics. We then apply this optimized model to elucidate the relaxation mechanisms responsible for the experimentally observed at G[10.1038/s41467-023-44494-3]. Finally, we extend our analysis to predict the dependence on the external magnetic field. These results offer valuable microscopic insight into the spin relaxation dynamics of V centers and establish a theoretical foundation for future developments in coherence control, defect engineering, and device optimization for quantum information applications.
@@ 2 Results
where the first term on the right-hand side of the equation describes the electron Zeeman interaction, while the second and third terms describe the zero-field splitting interaction. Here, is Bohr magneton, is an external magnetic field in the direction perpendicular to the hBN, is a factor of an electron, for an electronic spin-1, is the zero-field splitting equal to 3.479 GHz[10.1038/s41524-020-0305-x] at room temperature, and is the transverse component of the ZFS, which is set to zero in this work. We note that the unit of frequency (Hz), and not the angular frequency (rad-Hz), is used throughout this paper.
The electron spin Zeeman and zero-field splitting terms of are the leading terms in Eq. 1, which determine the main characteristics of the spectrum of the spin systems and have a major effect on the relaxation processes. At zero magnetic field, the eigenstates corresponding to and are split in the ground state due to the zero-field splitting. When an external magnetic field is applied, these eigenstates undergo additional splitting through the Zeeman interaction, resulting in the separation of the states, as shown in Fig. 1.
@@
where is the hyperfine-interaction tensor, calculated based on the method in Ref. [10.1038/s42005-024-01668-9] to tackle the finite-size effect. It is important to note that each branch in Fig. 1 is further divided into four sublevels due to hyperfine interactions with the three nuclear spins. These hyperfine splittings are approximately 66 MHz for 15N. At a magnetic field of approximately 1250 G, the eigenenergies of the and states approach near-degeneracy, see inset of Fig. 1. In this regime, strong mixing occurs between the electronic and nuclear spin states, enabling efficient spin flip-flop processes. This enhanced flip-flop interaction leads to pronounced correlations between spins, consequently modifying the T1 relaxation behavior, as will be discussed later.
As shown in Eq. (4), the , , , and hyperfine components govern the electron–nuclear spin flip-flop processes. To illustrate the strength of these interactions, their values are plotted in Fig. 2(b), highlighting the strongly coupled nature of the V defect system. Fig. 2(b) unravels that rotational invariant perpendicular hyperfine components of the first three nearest nitrogen nuclei are strong due to the Fermi contact contribution. This suggests that relaxation via the electron–nuclear spin flip-flop process in this channel is indispensable and plays a central role in the overall spin dynamics.
To simulate spin relaxation process, our models take into account all major parts of the spin Hamiltonian in Eq. (1) and truncate only the nuclear spin-nuclear spin dipolar couplings to a certain degree depending on the clustering used. The dimension of the given total Hamiltonian depends on the dimension of the Hilbert space of electron and nuclear spins, and the computational demand increases exponentially with the number of nuclear spins, limiting the capability of the exact solution. To resolve this issue, the methodology developed in Ref. [10.1103/PhysRevB.101.155203] is applied by dividing the entire system into smaller clusters of spins utilizing the central spin approximation and cluster expansion to reduce the Hilbert space considered while still maintaining all the aforementioned interactions.
@@
We simulate the dynamics of the system through the density matrix () following the below Master equation with an extended Lindbladian developed by Ref. [10.1103/PhysRevB.101.155203]
We note that represents an extended Lindbladian term introduced into the Master equation to account for spin flip-flop processes induced by environmental spins not included in cluster (diagonal elements in the density matrix), see Ref. [10.1103/PhysRevB.101.155203] for more details. Additionally, it is used to ensure the synchronization of key properties of different instances of the central spin across different clusters, such as expectation values. This correction is essential because the unsynchronization of a central spin in all clusters can lead to an artificial thermal state due to the accumulation of its magnetization. The extended term can be formulated as follows
Here, determines spin flip-flop rates of the central spin induced by environmental spins not incorporated in cluster and is the jump operator describing the spin flip and flop transitions of the central spin between and states. As such, Eq. (7) not only causes decoherence but also contributes to the decay of the central spin.
@@
2.1 Theoretical Prediction
As detailed in the Methods section, the time is determined from the population of the electron spin until it relaxes to thermal equilibrium. Fig. 1 illustrates the energy-level splitting among the three relevant states and suggests that the fields in the vicinity of the ground-state level anticrossing (GSLAC) at around 1250 G yield particularly narrow splitting, as the GSLAC is the point at which two ground-state sublevels become nearly degenerate. This implies that spin relaxation near GSLAC is faster, which can minimize computational expense. We therefore adopt a fixed field at 800 G to evaluate the performance of various computational models for predicting . This value is close to, but outside, the GSLAC regime.
2.2 Model 1: Two-Spin Clusters
@@
Fig. 4(a) depicts the evolution of the V electron spin initialized in the ms = 0 state. The result suggests that immediately after initialization, the electron spin polarization is rapidly transferred to the 15N nuclear spin in this model. At first, the relaxation appears exponential, but as the m population approaches the thermal equilibrium value of 1/3, as expected for the equal occupancy of the three spin sublevels (m), the relaxation pattern deviates from the exponential curve[10.1103/PhysRevB.65.205309, 10.1126/science.1155400] and exhibits a stretched exponential decay. This is confirmed by the exponent value of 0.37. Stretched exponents often arise when the relaxation is governed by multiple, independent decay mechanisms with different rates. While the distribution of the perpendicular hyperfine components, comprising strongly, moderately, and weakly coupled spins, see Fig. 2(b), could suggest such a multi-channel relaxation mechanism, experimental results do not support this model. As a consequence, we conclude that Model 1 fails to capture relevant correlation effects of the bath that may lead to different decay mechanisms.
2.3 Model 2 - Three/Four-Spin Clusters
@@
Unlike Model 1, this model incorporates dipolar interactions within each cluster, which may additionally contribute to spin dynamics behavior. In addition, Model 2 accounts for up to 4 spin correlation effects, thus, it significantly improves on Model 1 and can describe more complex dynamics. Here, 47 nitrogen clusters and 48 boron clusters have been used, describing a 238-spin model in total. Note that the number of clusters has been tested as shown in Supplementary S2.
@@
2.4 Model 3 - Three/Four-Spin Clusters with Additional Dephasing
Before moving toward another clusterization, we investigate the effect of decoherence of the electron spin on time. The cluster Hamiltonian in this model remains the same as in Model 2, and the number of clusters is also identical. However, in this model, we incorporate the effects of spin dephasing by introducing a dephasing time () of 200 ns [10.1038/s41467-023-44494-3] into the off-diagonal terms of the Lindbladian in Eq. (7). Note that the variation of values has been tested in Supplementary S3.
@@ 2.4 Model 3 - Three/Four-Spin Clusters with Additional Dephasing
2.5 Model 4 - Six-Spin Clusters
@@
To manage computational complexity, we apply the same reduced density matrix approach as in the previous two models. Furthermore, a total of 28 clusters are considered in this model, accounting for a total of 60 spins. It is important to note that in the model, we focus on nitrogen nuclear spins, and boron nuclei are neglected at first, as they exhibit a weak hyperfine interaction. Furthermore, previous studies have shown that heteronuclear spin flip-flop processes, e.g., 29Si-13C flip-flops in silicon carbide and 11B-15N flip-flops in hBN, are strongly suppressed at nonzero fields due to the differences in the gyromagnetic ratios of the nuclear spins.[10.1103/PhysRevB.90.241203] Consequently, in heterogeneous clusters consisting of the V electron spin, the three nitrogen nuclear spins in the central spin, and additional boron atoms, no direct coupling between the two types of nuclear spins is expected. We therefore assume that the nitrogen and boron nuclear baths decouple from each other, analogous to the case of decoherence.[10.1002/adfm.202511300] We note, however, that in contrast to SiC, the nuclear spins in hBN are much more strongly coupled to the electron spin, which may lead to electron-spin-mediated nuclear spin flip-flops[10.1002/adfm.202511300] and potentially to additional spin relaxation mechanisms. Therefore, the computed time obtained from Model 4 should be regarded as an upper bound. Note that the timestep convergence test and the effect of nuclear-spin synchronization have been validated in Supplementary S4.
@@
Model 4 enlarges the central‐spin subsystem to include the V electron spin and its three nearest‐neighbor 15N nuclei, while the bath is augmented by two additional distant 15N nuclei. The initial electron–spin polarization in the ms = 0 sublevel is set to 70%, in agreement with experimental measurements[10.1002/qute.202300118, 10.1103/PhysRevB.110.014104]. Note that the previous models initialized the electron spin polarization to 100% in the ms = 0 state for the sake of identifying the proper model. Since Model 4 includes all relevant interactions, we initialize the electron spin according to the experimental value of 70%. Under these conditions, the expected exponential relaxation of electron spin polarization appears, see Fig. 4(d). The decay behavior is well described by an exponential function (decay exponent equal to 1). This demonstrates that the previously unaccounted correlation effects arise from insufficient entanglement and dipolar interactions within the extended central spin and also with the bath spins. That is, the spin dynamics are mediated by the spin flip-flop processes. Specifically, referring to Eqs. (3) and (4), nuclear–nuclear flip–flop transitions are driven by the dipolar interaction term whereas electron–nuclear flip–flop transitions originate from the non‐secular hyperfine term . As illustrated in Fig. 2(b), the V electron spin exhibits strong hyperfine coupling to its three nearest nitrogen nuclei, therefore, including these four spins in the central‐spin description is essential to accurately capture the unseparable dynamics of the electron spin and the first neighbor nucelar spins. Simultaneously, the addition of two distant nuclear spins enables an accurate description of nuclear–nuclear flip‐flops via the dipolar interaction and additional correlation effects. Model 4 can be considered as defining the boundary between the central spin and the spin bath at the first-neighbor shell of the V center. Consequently, the reduced bath includes only the moderately and weakly coupled nuclear spins. In this extended central spin–moderately coupled spin bath model, the spin bath can be considered “memoryless”, or at least its correlations have no significant influence on the dynamics of the extended central spin system, which, in turn, gives rise to an exponential decay characteristic of Markovian processes. As a result, this model encompasses all significant relaxation pathways accurately and will be employed for all remaining calculations. A more detailed analysis of the contribution of distant nuclear spins is presented in the following section.